# PiCO: Peer Review in LLMs based on the Consistency Optimization

### v2

@@ 3.2 Performance Comparison
We validate the effectiveness of the proposed PiCO approach on three datasets by comparing the following two types of methods, i.e., the wisdom of the crowds and recent SOTA LLMs evaluation methods. The average results of PEN, CIN and LIS are demonstrated in Table 1. The ratios of response sets $\mathcal{D}$ are 1, 0.7, and 0.4, respectively.
Table 1 demonstrates that the proposed PiCO method consistently surpasses competing approaches across the majority of evaluated metrics. Notably, the PiCO approach achieves performance gains of 0.1, 2.5, and 0.92 on the PEN, CIN, and LIS metrics, respectively, compared to the Runner-up. These results underscore the superiority of aggregating evaluations from multiple models (such as Majority Voting, Rating Voting, PRD, and PRE) over relying on single-model methods like GPTScore and PandaLM. This collective model approach, leveraging ’the wisdom of the crowds’, more accurately aligns with human rankings in our open-question evaluation framework.
When comparing with PRD and PRE, which both incorporate peer review concepts, it is evident that the unsupervised PiCO method outperforms the supervised PRE approach. PRD, which adjusts model weights based on their win rates, shows promising results. Conversely, PRE distinctly relies on supervised human feedback data to conduct a qualification exam, subsequently assigning weights accordingly; however, it does not exceed the performance of the fully unsupervised PiCO method.
@@ 3.3 Exploring the Role of Confidence Weight
Figure 4: The heatmap distribution of the preference gap (PG) metric among seven LLMs across three datasets. A larger value (greater than 0) indicates a higher evaluation bias between two LLMs [17]. The first row represents the original PG value in three datasets, and the second row represents the re-weighting PG value by employing our learned confidence weight.
In this subsection, we will show that the confidence weight $w$ learned by our consistency optimization can reduce the system evaluation bias. Specifically, we first study whether the “review” model would prefer a particular model’s response. Following [17], we employ the preference gap (PG) to evaluate the bias as follows,
@@ 3.4 Study of Elimination Mechanism
Figure 5: Performance comparison of the PiCO (Ours) and PRE[17] methods on the Chatbot Arena, MT-Bench, and AlpacaEval datasets, with the number of eliminated reviewers on the x-axis. The y-axis is CIN, where lower values indicate better performance.
@@
The PiCO and PRE[17] methods both employ elimination mechanisms to remove those weakest LLMs from the “reviewer queue” during the evaluation process. To study the effect of the elimination mechanism, we plot the performance trend results in Figure 5. Here, the x-axis quantifies the number of reviewers eliminated, and the y-axis measures the CIN, where lower scores denote higher performance. Due to space limitations, more results on PEN and LIS metrics can be found in the Appendix E. It can be observed that both PiCO and PRE exhibit higher performance with an increasing number of eliminated “reviewers”. The proposed PiCO approach can achieve better performance than PRE in most cases. It is worth noting that the PRE method employs the accuracy of “qualification exams” to eliminate weak LLMs, and this process requires human annotation [17]. On the contrary, the elimination process of our PiCO method is unsupervised and can still achieve better evaluation results than PRE.
3.5 Validation of Consistency Assumption
In this subsection, we conduct the ablation study to validate the effectiveness of the consistency assumption, i.e., high-level LLM can evaluate more accurately than low-level ones while achieving higher scores. Specifically, we first manually construct three methods: Forward Weight Voting, Uniform Weight Voting, and Reverse Weight Voting. That is, the ability weights of the model are respectively weighted forward ($w=[1,0.9,...,0]$), uniformly ($w=[1,1,...,1]$), and backward ($w=[0,0.1,...,1]$) according to the ground-truth human ranking. Then, we randomly initialize the ability weights and employ our consistency optimization to adjust the weight. In addition, we also collect the average performance of “reviewer queue”, i.e., employing a single LLM as the “reviewer” to evaluate all response pairs and then calculate the average results of all LLMs.
The results are demonstrated in Table 2. It can be observed that the Forward Weight Voting achieves better results than the Uniform and Backward ones in all cases, while the Backward one achieves worse results. It validates that assigning larger weights to those models with stronger capabilities can obtain better results. Most importantly, employing our consistency optimization algorithm to assign weights to different review models can further improve the performance of the evaluation system, i.e., lower PEN and CIN, as well as higher LIS in all cases. Moreover, it is worth noting that the average performance of the “reviewer queue” is very poor, even worse than the Backward Weight Voting. This means that the answer-ranking data $\mathcal{D}$ contains a lot of evaluation noise, while the proposed approach can still optimize weights and obtain better ranking results. In summary, the above experimental results validate the effectiveness of the consistency assumption from various perspectives.
@@ 4 Related Work
With the proliferation of Large Language Models (LLMs), the task of effectively evaluating the generative capabilities of these LLMs has emerged as a formidable challenge. This section reviews the current mainstream methodologies for evaluating language models.
4.1 Evaluation Benchmarks for Diversity
LLMs are designed to tackle a broad spectrum of tasks, necessitating the development of benchmarks that can comprehensively assess their performance across various contexts[15]. GLUE[49] and SuperGLUE[48], with their focus on simulating real-world language processing scenarios, encompass a wide range of tasks including text classification, machine translation, reading comprehension, and dialogue generation. The HELM[34] benchmark diverges from task-specific assessments, offering a holistic evaluation of LLMs across dimensions such as language understanding, generation, coherence, context sensitivity, common sense reasoning, and domain-specific knowledge. The introduction of BIG-bench[43], featuring a diverse collection of 204 challenging tasks contributed by 450 authors from 132 institutions, aims to push the boundaries of language models’ capabilities. MMLU[28], covering 57 tasks across domains such as basic mathematics, American history, computer science, and law, seeks to measure the multitask accuracy of LLMs. Rankings based on performance in existing or novel evaluation benchmarks, such as OpenCompass[19] and C-Eval[29], have also been established. However, these evaluations may be compromised by the benchmark leakage issue, whereby the evaluation data is inadvertently used for model training, leading to an unjust enhancement of LLMs’ performance metrics[4; 33; 60].
Table 2: Ablation study comparing Backward, Uniform, Forward weight voting, and Consistency Optimization methods with the Average Performance of Reviewer Queue across three datasets, aiming for lower PEN and CIN scores to indicate better results.
Methods MT-Bench Chatbot Arena AlpacaEval
PEN ()($\downarrow$)( ↓ ) CIN()($\downarrow$)( ↓ ) PEN ()($\downarrow$)( ↓ ) CIN()($\downarrow$)( ↓ ) PEN ()($\downarrow$)( ↓ ) CIN()($\downarrow$)( ↓ )
Average Performance of Reviewer Queue $1.49^{\pm 0.28}$ $34.87^{\pm 14.68}$ $1.49^{\pm 0.26}$ $38.80^{\pm 19.28}$ $1.50^{\pm 0.23}$ $33.13^{\pm 13.97}$
Backward Weight Voting $1.43^{\pm 0.04}$ $25.00^{\pm 0.00}$ $1.43^{\pm 0.05}$ $26.00^{\pm 0.00}$ $1.36^{\pm 0.03}$ $24.00^{\pm 0.00}$
Uniform Weight Voting $1.34^{\pm 0.23}$ $22.00^{\pm 0.00}$ $1.39^{\pm 0.02}$ $24.00^{\pm 0.00}$ $1.34^{\pm 0.03}$ $22.00^{\pm 0.00}$
Forward Weight Voting $1.32^{\pm 0.03}$ $21.00^{\pm 0.00}$ $1.33^{\pm 0.03}$ $23.00^{\pm 0.00}$ $1.30^{\pm 0.05}$ $21.00^{\pm 0.00}$
Random Weight + Consistency Optimization $\bm{1.17^{\pm 0.06}}$ $\bm{17.50^{\pm 0.50}}$ $\bm{1.20^{\pm 0.08}}$ $\bm{18.00^{\pm 1.22}}$ $\bm{1.21^{\pm 0.04}}$ $\bm{19.00^{\pm 0.00}}$
4.2 Human Evaluation
@@
Human evaluation, aligning more closely with real-world application scenarios, offers effective and reliable feedback[15]. Through human assessment, Liang et al[34]. evaluated summary and misinformation scenarios across six models. Ziems et al[61]. engaged experts from diverse fields to assess model outputs across various domain datasets, aiding in the understanding of LLM performance in different CSS tasks. Bang et al[9]. assessed ChatGPT’s capabilities in Summarization, Machine Translation, and Reasoning Evaluations using datasets annotated by multiple humans. The LMSYS initiative introduced crowdsourced battle platforms like Chatbot Arena[59; 24], with human ratings serving as the primary evaluation metric. Despite the effectiveness of human annotation, its high cost and the potential for bias[17] and instability due to cultural and individual differences pose significant challenges[41].
4.3 Large Language Models for Evaluation
@@
The trend towards developing open-source LLMs has led to initiatives employing one or multiple LLMs as evaluators for assessing the outputs of LLMs. GPTScore[25], for instance, utilizes models like GPT-3 to assign higher probabilities to high-quality content through multidimensional evaluation prompted by multiple queries. Bubeck et al[12]. conducted a series of tests with GPT-4, finding its performance on multiple tasks to rival or even surpass human capabilities. Lin and Chen introduced the LLM-EVAL[35] methodology for automatically synthesizing evaluations of open-domain dialogue quality in large language models through single prompts. PandaLM[52] proposes a framework where an LLM acts as a "judge" for the automatic, reliable evaluation of LLMs tailored for instruction tuning optimization. G-EVAL[37] leverages GPT-4 and the Chain-of-Thoughts (CoT) paradigm, along with a tab-filling paradigm, for quality assessment of LLM outputs. Nevertheless, reliance on a single model for evaluation can introduce biases[31], such as positional bias[21], verbosity and length bias[50], and the favoring of the LLM’s own answers bias[37; 59]. In response, the ChatEval[14] framework proposes a multi-agent evaluation framework, simulating the human evaluative process with a multi-agent debate to enhance the automated assessment of text generation quality. Similarly, PRE[17] and PRD[31] have advocated for the use of LLMs as evaluators, amalgamating multiple evaluation outcomes for the automated assessment of other LLMs’ performance. However, like PRE method, which uses human feedback for supervised evaluation throughout the process, the comprehensive assessment of LLMs still incurs a relatively high cost.
@@ 5 Conclusion
In this paper, we propose PiCO, a novel LLM automatic evaluation direction without human feedback, utilizing peer-review mechanisms to measure LLMs automatically. In this setting, both open-source and closed-source LLMs lie in the same environment, capable of answering unlabeled questions and evaluating each other, where each LLM’s response score is jointly determined by other anonymous ones. Based on the consistency assumption, we maximize the consistency of each LLM’s capabilities and scores to re-rank the LLMs closer to human preferences. We propose three metrics called PEN, CIN, and LIS to measure the gap with human preferences. We perform experiments on multiple datasets and metrics, the results demonstrate that the proposed PiCO framework can effectively obtain an LLM leaderboard closer to human preferences. In the future, we will explore the peer-review mechanism for evaluating the ability of multi-modal large models.
Appendix A Dataset Format
@@
Appendix B Detailed Prompt for Reviewers

### v3

@@ 3.1 Performance Comparison
We validate the effectiveness of the proposed PiCO method on three datasets by comparing the following two types of methods, i.e., the wisdom of the crowds and recent SOTA LLMs evaluation methods. The average results with different rank-based metrics and datasets are demonstrated in Table 2. The ratios of response sets $\mathcal{D}$ are 1, 0.7, and 0.4, respectively.
The results presented in Table 2 demonstrate that the proposed PiCO method consistently outperforms competing approaches across most evaluated metrics, including surpassing all baselines, such as Claude-3 (API). Specifically, PiCO achieves improvements of 0.027, 0.047, and 0.14 on Spearman’s Rank Correlation Coefficient, Kendall’s Rank Correlation Coefficient, and Permutation Entropy metrics, respectively, compared to the runner-up. These results underscore the superiority of aggregating evaluations from multiple models, such as Majority Voting, Rating Voting, PRD, and PRE, as opposed to relying solely on single-model methods like GPTScore and PandaLM. This collective model approach, leveraging ’the wisdom of the crowds’, aligns with human rankings more accurately in our open-question evaluation framework.
In comparison with existing SOTA evaluation methods(i.e., PRD and PRE), it is evident that PiCO exhibits improvements across various evaluation metrics. Despite PRD’s adjustment of model weights based on their win rates and PRE’s reliance on supervised human feedback data to assign weights through a qualification exam, neither method achieves performance superior to the fully unsupervised PiCO approach. These methods rely on predefined criteria and human feedback, potentially leading to biases or suboptimal performance. In contrast, PiCO leverages unsupervised learning techniques, allowing it to autonomously adapt and discover patterns in the data without explicit human intervention.
@@
Figure 3: Heatmap distribution of preference gap (PG) metric among seven LLMs across three datasets. Higher values (above 0) indicate greater evaluation bias. The first row shows original PG values in three datasets, while the second row displays PG values re-weighted using our learned confidence weights.
3.2 Exploring the Role of Confidence Weight
In this subsection, we show that the confidence weight $w$ learned by our consistency optimization can reduce the system evaluation bias. Specifically, we first study whether the “review” model would prefer a particular model’s response. Following (Chu et al., 2024), we employ the preference gap (PG) to evaluate the bias as follows,
@@ 3.2 Exploring the Role of Confidence Weight
Figure 4: Performance comparison of the PiCO (Ours) and PRE methods on the Chatbot Arena, MT-Bench, and AlpacaEval datasets, with the number of eliminated reviewers on the x-axis. The y-axis is PEN, where lower values indicate better performance.
@@ 3.3 Study of Elimination Mechanism
Figure 5: The average loss for different numbers of eliminated reviewers($\downarrow$). It shows how the iterative elimination of weaker reviewers affects the overall loss in the peer-review system.
@@
Performance Comparison of Elimination Mechanisms. The PiCO and PRE methods both employ elimination mechanisms to remove those weakest LLMs from the “reviewer queue” during the evaluation process. As shown in Figure 4, the x-axis quantifies the number of reviewers eliminated, and the y-axis measures the PEN, where lower scores denote higher performance. It can be observed that both PiCO and PRE exhibit better performance with an increasing number of eliminated “reviewers”. The proposed PiCO approach can achieve better performance than PRE in most cases. It is worth noting that the PRE method employs the accuracy of “qualification exams” to eliminate weak LLMs, and this process requires human annotation (Chu et al., 2024). On the contrary, the elimination process of our PiCO method is unsupervised and can still achieve better evaluation results than PRE.
[…]
3.4 Other Results
Validation on more metrics (Precision@K and RBP@K). We demonstrated the results of precision and RBP (K=8,9,10) with other baselines in Table 3 (left). The results show that the proposed PiCO approach can achieve better precision and RBP performance in all cases. These results once again validate that PiCO can predict the LLM ranking more accurately than other baselines.
Comparison of tokens consumed. We compute the token consumption of each method in Table 3 (right). It can be observed that the proposed PiCO approach has a similar token consumed with other baselines (e.g., PRD and PRE) while achieving better evaluation performance. Although Chatbot Arena has a smaller token consumption, it requires 33k human annotations, while PiCO does not require any human annotations.
Stability validation of consistency optimization. We repeated the experiment with different seeds for 1000 times, and plotted the training loss curve and weight distribution in Figure 6. The results show that the proposed consistency optimization process is stable and the learned $w$ is convergence.
Comparing with existing benchmarks. We select the widely-used benchmarks (i.e., MMLU (Hendrycks et al., 2020) and GSM8K (Cobbe et al., 2021)) to evaluate the model performance ranking $\hat{\mathcal{R}}$, and calculate the Spearman’s $S(\uparrow)$ and Kendall’s $\tau(\uparrow)$ rank correlation with the human preference ranking $\mathcal{R}^{*}$. The results are demonstrated in Table 4. It can be observed that these benchmarks can only measure LLMs’ specific capability on a confined set of tasks, which fails to assess their alignment with human preference. These phenomena have been widely validated in other literature (Zhou et al., 2023; Zheng et al., 2023; Chang et al., 2023) and have almost become a consensus in the community of LLM evaluation.
Table 3: Comparison of more metrics (Precision@K and RBP@K) and token consumption on Chatbot Arena.
Methods RBP@K ()($\uparrow$)( ↑ ) Precision@K ()($\uparrow$)( ↑ ) Input Token Output Token Annotation Cost
8 9 10 8 9 10
Chatbot Arena Platforms - - - - - - $\sim 7500$k $\sim 10944$k $\sim 32$k
GPTScore(flan-t5-xxl) 26.2% 29.6% 45.1% 50.0% 55.6% 70.0% $\sim 22882$k $\sim 12260$k 0
GPTScore(davinci-002) 42.0% 50.6% 53.3% 62.5% 77.8% 80.0% $\sim 22882$k $\sim 12260$k 0
PandaLM 63.5% 63.5% 66.2% 62.5% 55.6% 60.0% $\sim 22882$k $\sim 10355$k 0
PRD 67.2% 73.8% 81.3% 87.5% 88.9% 80.0% $\sim 25087$k $\sim 10935$k 0
PRE 78.0% 81.3% 81.3% 87.5% 88.9% 80.0% $\sim 24120$k $\sim 11115$k $\sim 7$k
PiCO (Ours) 83.2% 83.2% 85.9% 100.0% 100.0% 90.0% $\sim 23823$k $\sim 11685$k 0
Figure 6: Stability validation of consistency optimization. We repeated the experiment with different seeds for 1000 times, and plotted the training loss curve and weight distribution. The results show that the learning process is stable and the learned $w$ is convergence.
@@ 4 Related Work
Evaluation Benchmarks for Diversity. LLMs are designed to handle a variety of tasks, necessitating comprehensive benchmarks (Chang et al., 2023). Notable benchmarks include GLUE and SuperGLUE (Wang et al., 2018; 2019), which simulate real-world scenarios across tasks such as text classification, translation, reading comprehension, and dialogue generation. HELM (Liang et al., 2022) provides a holistic evaluation of LLMs, assessing language understanding, generation, coherence, and reasoning. BIG-bench (Srivastava et al., 2022) pushes LLM capabilities with 204 diverse tasks. MMLU (Hendrycks et al., 2020) measures multitask accuracy across domains like mathematics and law. GSM8K (Cobbe et al., 2021) including 8K simple math questions with detailed solutions is widely used to assess the mathematical reasoning of models on grade-school-level questions. However, these evaluations can be compromised by benchmark leakage, where evaluation data inadvertently used for training leads to inflated performance metrics (Aiyappa et al., 2023).
Table 4: Comparison with existing benchmarks.
Benchmarks Spearman’s Rank Kendall’s Rank
Correlation Coefficient $S(\uparrow)$ Correlation Coefficient $\tau(\uparrow)$
MMLU 0.53 0.37
GSM8K 0.32 0.15
PiCO (Ours) 0.88 0.67
@@
Human Evaluation. Human evaluation provides reliable feedback that closely aligns with real-world applications (Chang et al., 2023). Liang et al. (2022) evaluated summary and misinformation scenarios across multiple models. Ziems et al. (2023) involved experts to assess model outputs in various domain-specific tasks. Bang et al. (2023) examined ChatGPT’s performance in summarization, translation, and reasoning using human-annotated datasets. The LMSYS initiative introduced platforms like Chatbot Arena (Zheng et al., 2023), relying on human ratings as the primary evaluation metric. Currently, using these anonymous battle platforms has become the primary way to evaluate the large language models, and its success is attributed to the wisdom of the crowds (Surowiecki, 2005; Budescu & Chen, 2015; Weller, 2007) and have been proven to lead to better results than that of an individual. Despite its effectiveness, human evaluation is costly and subject to bias and cultural differences(Peng et al., 1997).
@@
Large Language Models for Evaluation. The trend towards developing open-source LLMs has led to initiatives employing one or multiple LLMs as evaluators for assessing the outputs of LLMs. GPTScore (Fu et al., 2023) uses models like GPT-3 to assign probabilities to high-quality content through multidimensional evaluation. Bubeck et al. (2023) tested GPT-4, finding it rivaling human capabilities. Lin and Chen introduced LLM-EVAL (Lin & Chen, 2023) for evaluating dialogue quality with single prompts. PandaLM (Wang et al., 2023) employs LLMs as "judges" for evaluating instruction tuning. However, reliance on a single model can introduce biases such as positional (Dettmers et al., 2024), verbosity (Wang et al., 2024), and self-favoring biases (Liu et al., 2023b; Zheng et al., 2023). ChatEval (Chan et al., 2023) proposes a multi-agent framework to simulate human evaluation processes. Similarly, PRE (Chu et al., 2024) and PRD (Li et al., 2023a) use LLMs as evaluators, combining multiple evaluation outcomes for automated assessment. However, like the PRE method, which uses human feedback for supervised evaluation throughout the process, the comprehensive assessment of LLMs still incurs a relatively high cost.
@@ 5 Conclusion
In this paper, we propose PiCO, a novel unsupervised evaluation method to automatically evaluate Large Language Models (LLMs) without relying on human feedback. PiCO utilizes peer-review mechanisms to autonomously assess LLMs in a shared environment, where both open-source and closed-source models can respond to unlabeled questions and evaluate each other. In this setup, each LLM’s response score is determined collectively by other anonymous models, aiming to maximize consistency across capabilities and scores. The extensive experiment results across multiple datasets and standard rank-based metrics demonstrate that PiCO effectively generates an LLM ranking that aligns closely with human preferences. In the future, we plan to extend the peer-review mechanism to evaluate the capabilities of multi-modality large models.
Acknowledgements
This work was supported in part by the Natural Science Foundation of China (No. 62202014, 62332002, 62425101, 62088102), and also supported by the China Postdoctoral Science Foundation under Grant Number BX20240013 and 2024M760113.
Appendix A Detailed Explanation of Metrics
In this section, we provide a comprehensive explanation of the metrics used to evaluate the alignment between learned LLM rankings and human rankings. These metrics assess the strength of correlations, complexity, and the level of agreement between rankings. Specifically, we discuss five key metrics: Spearman’s Rank Correlation Coefficient, Kendall’s Rank Correlation Coefficient, Permutation Entropy, Count Inversions, and Longest Increasing Subsequence, detailing their formulations and intuitive interpretations.
i) Spearman’s Rank Correlation Coefficient $S(\uparrow)$ (Lehman et al., 2013) measures the strength and direction of the monotonic relationship between two ranked variables. It is computed as:
$S(\hat{\mathcal{R}},\mathcal{R}^{*})=1-\frac{6\sum_{i=1}^{m}d_{i}^{2}}{m(m^{2}-1)},$ (11)
where $d_{i}=\text{rank}_{\hat{\mathcal{R}}}(M_{i})-\text{rank}_{\mathcal{R}^{*}}(M_{i})$ is the difference between the ranks of LLM $M_{i}$ in the learned ranking $\hat{\mathcal{R}}$ and the human ranking $\mathcal{R}^{*}$, and $m$ is the total number of LLMs. A higher Spearman coefficient indicates a stronger correlation between the rankings.
ii) Kendall’s Rank Correlation Coefficient $\tau(\uparrow)$ (Kendall, 1938) evaluates the similarity between two rankings by counting the number of concordant and discordant pairs. It is given by:
$\tau(\hat{\mathcal{R}},\mathcal{R}^{*})=\frac{C-D}{\frac{1}{2}m(m-1)},$ (12)
where $C$ represents the number of concordant pairs, and $D$ represents the number of discordant pairs. A pair $(M_{i},M_{j})$ is concordant if $M_{i}$ and $M_{j}$ have the same order in both $\hat{\mathcal{R}}$ and $\mathcal{R}^{*}$, meaning if $M_{i}\succ M_{j}$ in $\hat{\mathcal{R}}$, then $M_{i}\succ M_{j}$ in $\mathcal{R}^{*}$. Conversely, a pair is discordant if their relative order differs between the two rankings. A higher $\tau$ value indicates a closer alignment between the rankings.
iii) Permutation Entropy $H(\downarrow)$ (Bandt & Pompe, 2002) measures the complexity or randomness of sequences, which is formulated as follows:
$H(\hat{\mathcal{R}},\mathcal{R}^{*}):=-\sum p(\pi)\log p(\pi),$ (13)
where
iv) Count Inversions $C(\downarrow)$. Counting inversions (Leiserson et al., 1994) aims to measure the degree of disorder or "invertedness" in an array or sequence of elements. We thus define it as follows,
$C(\hat{\mathcal{R}},\mathcal{R}^{*}):=\sum_{M_{i},M_{j}\sim\mathcal{M}}\mathbf{1}\{M_{i}\succ M_{j}\land i<j\}.$ (14)
v) Longest Increasing Subsequence $L(\uparrow)$. The longest increasing subsequence aims to find the length of the longest subsequence in a given sequence of elements, where the subsequence is in increasing order. We utilize it to measure the degree of match with human rankings as follows,
$L(\hat{\mathcal{R}},\mathcal{R}^{*}):=\max\left\{dp[i]\mid 1\leq i\leq m\right\},$ (15)
where
$dp[i]=1+\max\left\{dp[j]\mid 1\leq j<i\land M_{j}\prec M_{i}\right\}.$
Appendix B Dataset Format
@@
Appendix C Detailed Prompt for Reviewers